flying with the trail wind an airplane flies 600 km per hour. Flying with the same wind it flies 500 km per hour
to make the return flight against the wind. Find the speed of the wind and speed of the airplane.

Answer by Alan3354(69443)

The airspeed of the plane is the average = (600+500)/2 = 550
The windspeed is the difference = 600 - 550 = 50
----------------
PS Airplanes don't use km/hr, they use knots.
1 knot = 1 nautical mile per hour.